OpenStudy (anonymous):

Find the solution to the equation 3(y - 5) = -12.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

3(y-5)=-12 3y-15=-12 +15 +15 3y=3 - - 3 3 y=1

OpenStudy (owlcoffee):

\[3(y-5)=-12\] First off, you can begin by applying distributive property to simplify the expression and make it into more simpler terms: \[3y-15=-12\] Then isolate the 3y: \[3y=-12+15\] which simplifying is: \[3y=3\]

OpenStudy (anonymous):

You're supposed to get y alone so you divide the three off the y because its multiplying but you have to do what you did to one side to the other side, and 3 divided by 3 is 1.